OpenStudy (anonymous):

Rectangles A and B are similar. The area of A is 30.195 cm^2. The length of B is 6.1 cm. Each dimension of B is 2/3 the corresponding dimension of A. What is the perimeter of B?

OpenStudy (mathstudent55):

Each dimension of B is 2/3 the corresponding dimension of A. That means (2/3)A = B If you solve that for B, you get: A = (3/2)B which is the same as A = 1.5B That means every dimension in A is one-and-a-half times the corresponding dimension in B. Since you are given that the length of B is 6.1 cm, if you multiply that by 1.5, you get the length of A. Then use the area of A and divide by the length of A to find the width of A. Then to find the perimeter of A, add the length + length + width + width of A.